Suppose we have a market demand Q = 18 – P and a cost C(Q) 9) = 3Q?.Suppose that that firm 2 that invests in a new technology that changes it cost structure from firm 1. Market demand is still
Answer #1

as For firmie Q = 18-1 p=18-0 Here, Q = 9,792 MC, dCTC ;) dq, q IMC, 2 The = Pixh (18-9-219, = 189, - q, 2-9q TRE de TRJ dh aand price, p= 20-20 = 20-2x 4 20-8 $ 12 MC₂ = a (T(2) dqz fx 292 . 92/3 i. R2 = 1892 - 99₂ - 9 2² FR2 & CTR 22 dq2 a 18- q -c> Now folving for q , 42 using elimination Method 39, +92 - 18 39, +792 = 54 692= 36 96 and 39,7 q2 = 18 G = 18-6 39 = 12 9,